### v1

@@ Abstract
This paper considers the need for generalizable bias mitigation techniques in machine learning due to the growing concerns of fairness and discrimination in data-driven decision-making procedures across a range of industries. While many existing methods for mitigating bias in machine learning have succeeded in specific cases, they often lack generalizability and cannot be easily applied to different data types or models. Additionally, the trade-off between accuracy and fairness remains a fundamental tension in the field. To address these issues, we propose a bias mitigation method based on multi-task learning, utilizing the concept of Monte-Carlo dropout and Pareto optimality from multi-objective optimization. This method optimizes accuracy and fairness while improving the model’s explainability without using sensitive information. We test this method on three datasets from different domains and show how it can deliver the most desired trade-off between model fairness and performance. This allows for tuning in specific domains where one metric may be more important than another. With the framework we introduce in this paper, we aim to enhance the fairness-performance trade-off and offer a solution to bias mitigation methods’ generalizability issues in machine learning.
Keywords Bias  $\cdot$ Fairness  $\cdot$ Monte-Carlo dropout  $\cdot$ Pareto Optimality
@@ 1 Introduction
The term bias was first introduced in the machine learning domain by Tom Mitchell in his 1980 paper titled “The need for biases in learning generalizations” Mitchell (1980). The concept of bias refers to giving importance to particular features to improve generalization. This general idea of bias in machine learning is positive and necessary for models to perform, eliminating the risk of hyper-focusing on specific samples over others. On the contrary, bias can also be negative in machine learning. Negative bias can be defined as an inaccurate assumption made by a machine learning algorithm that is systematically or historically prejudiced against certain groups of people Zanna et al. (2022). Decisions made by these biased algorithms could cause adverse effects on particular social groups, for example, those defined by sex, race, age, marital status, handicaps, etc., when used to make autonomous decisions in life-changing cases such as health, hiring, education, criminal sentencing, etc.
Negative bias can be introduced into the machine pipeline in two main ways, through the data or the algorithm itself Blanzeisky and Cunningham (2021). Bias due to data, also known as a negative legacy Cunningham and Delany (2021); Kamishima et al. (2012), can be caused by an imbalance in the representation of different population categories in the training data, meaning if certain demographics are lacking from the sample data, models trained on this data often do not generalize when applied to new data that contains those missing demographics Gorrostieta et al. (2019). Another cause of negative legacy is labeling errors. This can occur because data fed into models, especially supervised or semi-supervised models which are widely used in various jurisdictions, are manually labeled by humans who are inherently biased. These models can ultimately reflect people’s impressions and sustain or further magnify bias from the labeled data Vallor (2017). Bias caused by the algorithm itself is regarded as underestimation, and it occurs when the classifier under-predicts an uncommon outcome for the minority group. This can happen when the classifier under-fits the data, causing it to focus on strong signals and missing more subtle phenomena Cunningham and Delany (2021).
For over 10 years, research on bias and fairness in machine learning has been extensive and numerous works have explored the development of analytical techniques that are aware of social fairness and discrimination Kamishima et al. (2012); Friedler et al. (2019). These studies have shown that simply eliminating the sensitive features from the model’s workings is insufficient to avoid an unfavorable determination process. This is due to the indirect influence of sensitive information on other features vital to the algorithm’s decision-making process. This highlights the need and emergence of bias mitigation techniques.
Despite their seeming success in specific cases, there is a recurring trend of bias mitigation methods lacking generalizability. With applications of machine learning in data-driven decision-making procedures in a rapidly growing number of industries and the fairness concerns that arise from them, there is an increasing need for generalizable and more easily transferable bias mitigation techniques. This is because the machine learning algorithms developed and deployed on large scales are usually distinct and often proprietary and have different concerns with bias and limitations when choosing methods to mitigate it. For instance, searches for distinctively Black-sounding names are more likely to trigger ads for arrest records Sweeney (2013). In contrast, image searches for professions such as CEO produce fewer images of women Obermeyer et al. (2019), and natural language processing algorithms encode language in gendered ways Caliskan et al. (2017), among other examples. These are all highly different types of data and models that often cannot be analyzed or manipulated in the same manner. Unfortunately, most existing solutions for mitigating bias and unfairness are one-off applications for a specific problem or use case Pagano et al. (2022). This calls for a need to develop more generalizable bias mitigation techniques that can be implemented with various model types.
@@
Another setback with developing methods to mitigate bias that researchers face is the trade-off between accuracy and fairness, which is prevalent in existing literature Dutta et al. (2020). This challenge proves the need for methods that reconcile this fundamental tension in the bias and fairness aspect of machine learning. As a solution to this issue, several previous works have used the concept of Pareto optimality from multi-objective optimization to seek the fairness-accuracy Pareto front of a neural network classifier Wei and Niethammer (2022); Liu and Vicente (2022); Wang et al. (2021a); Liang et al. (2022); Little et al. (2022); Kamani et al. (2021). Before the introduction of this concept to fairness literature within the past 2 years, existing algorithmic fairness methods have typically performed the so-called linear scalarization scheme, which has limitations in recovering Pareto optimal solutions Wei and Niethammer (2022).
In this work, we propose a bias mitigation method based on multi-task learning (MTL), which has been used in previous research to optimize both accuracy and fairness Oneto et al. (2019), along with Monte Carlo dropout to utilize model uncertainty to improve fairness without a massive sacrifice in either computational complexity or performance. Our method is an extension of our previous work Zanna et al. (2022), a hybrid in-processing and post-processing technique that does not require access to a machine learning model’s objective function and prediction methodology. This method enables us to optimize accuracy and fairness while improving the explainability of the model without explicitly using any of the sensitive information in the functional form of our final model. Researchers have mentioned that analyzing bias and unfairness collaborates with model explainability, so explainability is intrinsic to the study Pagano et al. (2022). We also utilize the concept of Pareto optimality from multi-objective optimization to arrive at models that deliver the most desired trade-off between fairness and performance. This allows the results to be tuned according to the needs in any domain where one metric may be more important than another. We analyze and test our method on two publicly available (ADULT & MIMIC-III) and 1 non-public (SNAPSHOT) dataset from three varying domains: in-hospital mortality, finance, and stress prediction, to demonstrate how transferable the method is between different fields. One such dataset is one of the most well-used data for bias studies (ADULT dataset).
@@ 2 Related Works
2.1 Definition of Fairness
@@
The concept of fairness has been around in philosophy and psychology and, recently, in machine learning. Current literature has shown a consensus that fairness has to be studied as the endpoint of implementing algorithms Pleiss et al. (2017). Enforcing fairness when building machine learning models requires that fairness is first defined. Philosophy and psychology have tried to define the concept of fairness long before computer science, and the fact that there is no universally agreed-upon definition of fairness further complicates the problem Saleiro et al. (2018). Generally, fairness can be defined as the absence of prejudice or favoritism towards an individual or a group based on their intrinsic or acquired traits in decision-making Saxena et al. (2019). Even though the necessity of fairness has been demonstrated through numerous perspectives, it can be challenging to achieve in practice in machine learning and society as a whole Mehrabi et al. (2021). In this section, we explore different classes of bias mitigation techniques used to attempt to achieve different definitions of fairness on data from different domains and how effective they are for these cases.
2.2 Bias Mitigation Techniques
When it comes to methods developed to counter the issue of negative bias, there are several techniques Calders and Verwer (2010); Chouldechova (2017); Feldman et al. (2015); Hardt et al. (2016); Kamiran and Calders (2012); Zafar et al. (2017); Zhang et al. (2018) that are generally categorized into 3 groups: pre-processing, in-processing, and post-processing methods. Most of these existing bias mitigation methods try to avoid unethical interference of sensitive (protected) attributes in the functional form of models, while others try to include participants from disadvantaged groups Mehrabi et al. (2021). In addition, some works try to satisfy one or more of the fairness notions in their methods, such as disparate impact (DI) ratios, differences in false negative and true positive scores, etc., to satisfy notions of treatment disparity and impact disparity by allowing the protected attributes during the training phase, but avoiding them at prediction Mehrabi et al. (2021). This section will discuss some of these techniques from recent literature, their weaknesses, and what our proposed method provides to bridge the gaps.
2.2.1 Pre-processing Techniques
Perhaps the easiest bias mitigation methods to develop and implement are pre-processing techniques. These involve eliminating the sensitive attributes, changing the class labels, and reweighing or resampling data to remove imbalances without relabeling instances Kamiran and Calders (2012). Pre-processing techniques are used if the algorithm is allowed to modify the training data. They are generally easier to implement, and there is no need to access sensitive labels at test time, which is not allowed in some jurisdictions d’Alessandro et al. (2017); Yan et al. (2020). A good example of this class of techniques is by the authors of Bryant et al. (2019), who implement a pre-processing method to remove sensitive data (e.g., race, gender, etc.) from the model for a financial banking system. They aimed to ensure customer data protection, anonymization, and fairness without affecting the machine learning model. To achieve this, they generate synthetic data from the representation of the original data in order to preserve privacy while maintaining statistical similarity to the original data using Variational Auto-Encoders (VAE). Their method performed well on some datasets they tested. Another work that utilized the method of data synthesizing is Jang et al. (2021), where the authors generated a new dataset that does not contain disparity of distribution, quality, or noise to ensure that all classes are treated equally by building the model on purely generated data that is balanced (to ensure fairness), and then transferring to real data for fine-tuning (to ensure performance). They also utilized the VAE-GAN to approximate the data distribution and generate data with desired sensitive attributes and labels. Georgopoulos et al. (2021) also utilized the idea of synthesizing data in computer vision. A bias mitigation method was proposed to increase demographic diversity in a facial image dataset using Generative Adversarial Networks (GANs) to create additional images by transferring demographic attributes to each image in the unbalanced dataset. These methods generally improve fairness, but they are very domain-oriented, and the same technique cannot be implemented on various data types. There is also some domain-related limitation where for example with health-related data, generation of synthetic data can be tricky. Chen et al. (2021) states that the roles of synthetic data in AI systems in healthcare remain unclear. Specifically, when generating synthetic medical data for rare or new disease conditions, there may not be sufficient samples to establish clinical reference standards. Some works utilize the reweighing pre-processing technique for bias mitigation Paviglianiti and Pasero (2020); Vega-Gonzalo and Christidis (2022); Krasanakis et al. (2018). This method has shown to work reasonably well in improving fairness, but it usually struggles when the dataset does not contain sufficient samples for a specific demographic group Yang et al. (2022).
Our proposed method which is a hybrid between in-processing and post-processing does not incur the domain-related limitations that pre-processing techniques typically suffer from.
2.2.2 In-processing Techniques
In-processing techniques try to modify and change learning algorithms to remove discrimination during the model training process Gorrostieta et al. (2019) - either by incorporating changes into the objective function or imposing a constraint Liu and Vicente (2022). If it is allowed to change the learning procedure for a machine learning model and the sensitive attributes information is accessible in the training samples, in-processing techniques can be used during the training of a model Yan et al. (2020); Gorrostieta et al. (2019); Oneto et al. (2019). In recent years, in-processing techniques have become more prevalent Pagano et al. (2022) due to them generally resulting in better performance on accuracy and fairness. For example, Radovanović et al. (2020) presents a bias mitigation technique targeting group bias. In this work, they modify the logistic regression algorithm to add fairness constraints into the model learning process, forcing the predictive model not to create disparate impact and allow equal opportunity for every sub-population. The authors also reiterated the advantage of in-processing techniques over pre-processing since they do not require altering the data, which hurts ethical norms.
Another in-processing method is introduced in Ashokan and Haas (2021) where a mitigation strategy in which the algorithm learns the difference between predicted and observed ratings in subgroups. The algorithm identifies which is biased and corrects the prediction. Fairness here improved in most use cases, but performance highly varied. Du et al. (2021) attempts to make Deep Neural Networks (DNNs) fairer by targeting the bias only in the output layer. The authors manipulate the structure of the model and regularize its training, resulting in training samples with different sensitive attributes being neutralized. This method reduced the model’s dependence on the sensitive attribute and subsequently improved fairness. This method performed well in terms of fairness but came with a high performance loss.
Despite these aforementioned methods seemingly delivering all they need in terms of fairness, in-processing techniques are also task-specific since they involve manipulating the model itself, which makes them non-transferable, and sometimes modifying the model may not be possible Oneto et al. (2019). In addition to this, these methods often come at a high performance cost. With our proposed method, we tackle the high performance loss by utilizing the Pareto optimality concept that allows models to be tuned to a particular domain where one metric, either fairness or performance, may be more important than the other. Our method also does not require access to a machine learning model’s objective function and prediction methodology, eliminating the issue of in-processing techniques having to be task-specific or modifying the model.
2.2.3 Post-processing Techniques
These methods aim to edit posteriors in ways that satisfy the fairness constraint, and no changes are made to the model Oneto et al. (2019). Post-processing techniques can be used in a scenario where a machine learning algorithm can only treat the learned model as a black box without any ability to modify the training data or learning algorithm. The advantages of this class of techniques are that they can be applied to any classifier since they eliminate the need for modifications. They also have relatively good performance, especially regarding fairness measures.
As a recent form of post-processing for bias mitigation in models, several prior works Zhang et al. (2018); Li et al. (2022); Grari et al. (2019); Adel et al. (2019) have utilized adversarial learning to mitigate bias. Adversarial learning is a method that aims to trick machine learning models by providing deceptive input Huang et al. (2011). It includes generating and detecting adversarial examples, which are inputs specially created to deceive classifiers. Zhang et al. (2018) presents a framework for mitigating bias by including a variable for the interest group and simultaneously learning a predictor and an adversary. The authors claimed that in theory, the method could enforce the chosen metric across multiple definitions of fairness, regardless of the complexity of the predictor’s model. As a result, they were able to train models that were significantly less biased but could not get these models to converge in the end. As for Li et al. (2022), their work focused on studying the fairness of outlier detection. They propose an outlier detection method that utilizes adversarial networks to learn the optimal representation of the original data while hiding the sensitive attribute in the data. Their method was able to improve group fairness and outlier detection performance. In Grari et al. (2019), Grari et al. developed an approach of adversarial gradient tree boosting where the objective is to predict the output with gradient tree boosting while minimizing the ability of an adversarial neural network to predict the sensitive attribute. This approach achieved a similar level of fairness to other proposed algorithms. And lastly in Adel et al. (2019), the authors made slight adjustments to their architecture, given any existing differentiable classifier by adding a new hidden layer, to enable concurrent adversarial optimization for fairness and accuracy.
The general drawback to post-processing methods is that they do not allow for flexibility with the accuracy-fairness trade-off Oneto et al. (2019); Yan et al. (2020); Gorrostieta et al. (2019), which we aim to provide with our proposed method. Additionally, with the current bias mitigation methods, there exists minimal support for black-box models compared to the abundance of information for white-box models. There is also a need for transparency and explainability in machine learning models and mitigation methods that are transferable across various model structures and data types. With our work, we aim to address some of these concerns.
3 Methods
3.1 Fairness Terminology

### v2

@@ Abstract
Bias originates from both data and algorithmic design, often exacerbated by traditional fairness methods that fail to address the subtle impacts of protected attributes. This study introduces an approach to mitigate bias in machine learning by leveraging model uncertainty. Our approach utilizes a multi-task learning (MTL) framework combined with Monte Carlo (MC) Dropout to assess and mitigate uncertainty in predictions related to protected labels. By incorporating MC Dropout, our framework quantifies prediction uncertainty, which is crucial in areas with vague decision boundaries, thereby enhancing model fairness. Our methodology integrates multi-objective learning through pareto-optimality to balance fairness and performance across various applications. We demonstrate the effectiveness and transferability of our approach across multiple datasets and enhance model explainability through saliency maps to interpret how input features influence predictions, thereby enhancing the interpretability of machine learning models in practical applications.
@@ 1 Introduction
Introduced by Tom Mitchell in 1980 (Mitchell (1980)), the term "bias" in machine learning refers to giving importance to particular features to improve generalization—which is essential for model performance. On the contrary, bias can also be negative, leading to inaccurate assumptions by the algorithm that can be prejudiced against certain groups of people (Zanna et al. (2022)). These decisions could cause adverse effects on particular social groups, for example, those defined by sex, race, age, marital status, handicaps, etc., when used to make autonomous decisions in life-changing cases such as health, hiring, education, criminal sentencing, etc.
Negative bias arises either from the data itself or the algorithm’s design (Blanzeisky and Cunningham (2021)). Data-related biases, or ’negative legacies’ (Kamishima et al. (2012)), stem from unrepresentative training samples or labeling errors, while algorithmic biases or "underestimation" occurs when the classifier under-predicts an uncommon outcome for the minority group. This can happen when the classifier under-fits the data, causing it to focus on strong signals and missing more subtle phenomena (Cunningham and Delany (2021)).
In recent years, research on bias and fairness in machine learning has shown that simply eliminating the sensitive features from the model’s workings is insufficient to avoid an unfavorable determination process. This is due to the indirect influence of sensitive information on other features vital to the algorithm’s decision-making process. As mentioned by Paulus and Kent (2020), the use of “high dimensional” or “black box” prediction techniques is generally problematic given that these approaches can predict the protected label through other variables, whether or not the actual protected label is explicitly encoded. Given these scenarios, it is understood that reducing model bias may not be sufficient in eliminating fairness concerns in decision contexts associated with certain fields such as predictive healthcare, which can be complex, unintuitive, and often hard to explain (Yang (2022)). This highlights the need for bias solutions to strive for better explainability to improve understanding of the model, increase the utility of its output, and produce better outcomes (Yang (2022)).
@@
Another setback with developing methods to mitigate bias is the trade-off between performance and fairness, which is prevalent in existing literature (Pessach and Shmueli (2022); Buijsman (2023)). As a solution to this issue, several previous works have used the concept of Pareto optimality from multi-objective optimization to seek the fairness-accuracy Pareto front of a neural network classifier (Wei and Niethammer (2022); Liu and Vicente (2022); Wang et al. (2021a); Liang et al. (2022); Little et al. (2022); Kamani et al. (2021)). Multi-objective optimization addresses the problem of optimizing a set of possibly contrasting objectives (Sener and Koltun (2018)). Before the introduction of this concept to fairness literature within the past 4 years, existing algorithmic fairness methods have typically performed the so-called linear scalarization scheme. This is a multi-objective optimization method that converts multiple objectives into a single composite objective (Marler and Arora (2004)). It can miss significant trade-offs between objectives because it assumes a linear relationship among them, which presents limitations in recovering Pareto optimal solutions (Wei and Niethammer (2022)).
As machine learning plays a more significant role in decision-making across industries, the demand for domain-agnostic and easily transferable bias mitigation techniques is increasing. Many machine learning algorithms are unique and proprietary, each facing different challenges related to bias. Current mitigation solutions often target specific problems and are not flexible enough to address varying fairness and performance concerns (Pagano et al. (2022)). This situation highlights the need for versatile bias mitigation methods suitable for diverse domains.
Our research introduces a bias mitigation approach that utilizes model uncertainty to enhance fairness in machine learning. Our method leverages epistemic uncertainty which arises from incomplete knowledge about model parameters and can be reduced with additional data (Stone et al. (2022)). This uncertainty correlates with model weights and highlights biases which can be addressed through mitigation strategies (Li et al. (2021)). Specifically, using weights from models uncertain about a protected label helps diminish its influence on the outcome by averaging over multiple plausible models.
To evaluate the uncertainty of the protected label, we employ a multi-task learning (MTL) framework to predict target and protected labels while assessing their associated uncertainties. This framework enhances learning by sharing information across tasks, improving model efficacy (Caruana (1997)). To model uncertainty, we employ Monte Carlo (MC) Dropout to perform multiple forward passes with varied dropout masks, creating a distribution of outputs to estimate uncertainty accurately (Milanés-Hermosilla et al. (2021)). The ensemble approach of MC Dropout during inference maintains high performance by reducing the impact of random dropout noise in individual predictions (Ahmed et al. (2023)). This technique also quantifies prediction uncertainty, essential in regions with vague decision boundaries or insufficient data (Choubineh et al. (2023)), which often disproportionately affects marginalized groups, underscoring the importance of fairness (Kendall and Gal (2017)). By focusing adjustments on areas of highest uncertainty, especially in sensitive applications, our method enhances fairness and model robustness.
Extending our prior work (Zanna et al. (2022)), we also integrate multi-objective learning via pareto-optimality into this framework, allowing us to develop models that achieve the optimal balance between fairness and performance. This flexibility ensures that the results can be tailored to meet specific needs across various domains, where one metric might be prioritized over others. We further analyze and test this framework on multiple classification-based datasets, including two widely-used public datasets (ADULT & MIMIC-III) and a private dataset (SNAPSHOT) across diverse domains such as in-hospital mortality, finance, and stress prediction. This demonstrates the method’s transferability across different fields. Notably, the ADULT dataset is extensively utilized in bias studies, highlighting its significance in this research.
Furthermore, our approach enhances model explainability without incorporating sensitive data directly. By employing a saliency map technique (Simonyan et al. (2013)), we interpret how input features influence predictions and highlight our method’s impact on feature prioritization. This feature will allow practitioners in various fields to interpret better the decision-making process of the models they utilize. The major contributions of this work include:
  • Introduced a bias mitigation approach that leverages MTL and MC Dropout which operates independently of the machine learning model’s objective function and prediction methodology, avoiding the use of sensitive data in the model’s functional form.
  • Integrated Multi-Objective Learning Through Pareto-Optimality to balance fairness and performance, tailoring outcomes to specific domain needs where certain metrics may be prioritized.
  • Validated the proposed method on several classification-based datasets, including the ADULT and MIMIC-III public datasets, and a non-public SNAPSHOT dataset from three domains.
  • Compared performance of the proposed method to other baseline methods namely: Reweighting, ARL, and FairRF.
  • Employed a saliency map technique to visualize the significance of model weights, improving our understanding of how input features influence final predictions. This visualization also highlights the impact of our proposed method on the network’s prioritization of features for prediction.
@@ 2 Related Works
2.1 Fairness Challenges in Machine Learning
@@
Fairness has long been studied in philosophy and psychology and has recently become critical in machine learning (Pleiss et al. (2017)). Generally, fairness can be defined as the absence of prejudice or favoritism towards an individual or a group based on their intrinsic or acquired traits in decision-making (Saxena et al. (2019)). Even though the necessity of fairness has been demonstrated through numerous perspectives, it can be challenging to achieve in practice in machine learning and society as a whole (Mehrabi et al. (2021)). There are several bias mitigation techniques (Calders and Verwer (2010); Chouldechova (2017); Kamiran and Calders (2012); Zafar et al. (2017); Zhang et al. (2018)) introduced in prior literature, which are generally categorized into 3 groups: pre-processing, in-processing, and post-processing methods.
Pre-processing techniques are used if the algorithm is allowed to modify the training data. They are generally easier to implement, and there is no need to access sensitive labels at test time, which is not allowed in some jurisdictions (d’Alessandro et al. (2017); Yan et al. (2020)). These methods improve fairness, but they are very domain-oriented and could pose some domain-related limitations where for example with health-related data, generation of synthetic data can be tricky (Chen et al. (2021)). Some works utilize the reweighing pre-processing technique for bias mitigation (Paviglianiti and Pasero (2020); Vega-Gonzalo and Christidis (2022); Krasanakis et al. (2018)). This method has shown to work reasonably well in improving fairness and has become a standard in the field, but it usually struggles when the dataset does not contain sufficient samples for a specific demographic group (Yang et al. (2022)). Our proposed method which is a hybrid between in-processing and post-processing does not incur the domain-related limitations that pre-processing techniques typically suffer from.
In-processing techniques modify learning algorithms to eliminate discrimination during model training by altering the objective function or adding constraints (Gorrostieta et al. (2019); Liu and Vicente (2022)). These techniques are applicable when changes to the learning process are permissible, and sensitive attribute data is available (Yan et al. (2020); Gorrostieta et al. (2019)). In recent years, in-processing techniques have gained popularity due to their enhanced accuracy and fairness (Pagano et al. (2022)). However, these methods are typically task-specific and non-transferable because they require direct modifications to the model, which may not always be feasible and can incur high performance costs (Oneto et al. (2019)). Our proposed method overcomes these limitations by utilizing Pareto optimality, allowing for flexible tuning between fairness and performance. Additionally, it does not necessitate changes to the machine learning model’s core functions, thus eliminating the need for task-specific alterations commonly associated with in-processing techniques.
Post-processing bias mitigation methods aim to edit posteriors in ways that satisfy the fairness constraint, and no changes are made to the model (Oneto et al. (2019)). These techniques are suitable when a machine learning model must be treated as a black box, without modifications to training data or algorithms. They are applicable to any classifier and generally perform well in fairness. However, they lack flexibility in balancing accuracy with fairness (Oneto et al. (2019); Yan et al. (2020)), an issue our method addresses using Pareto optimality. Current bias mitigation methods also offer limited support for black-box models, unlike the extensive resources available for white-box models.
2.2 Multi-task Learning for Bias Mitigation
Multi-task learning is a machine learning technique where a model is trained to perform multiple tasks simultaneously. It aims to leverage useful information contained in multiple related tasks to help improve the generalization performance of all the tasks (Zhang and Yang (2018)). By doing so, MTL not only optimizes the accuracy of each task but also enhances fairness by ensuring that the model does not disproportionately learn from or favor any single task or demographic group. Studies have demonstrated that models trained using MTL not only perform better on average but also show improved fairness metrics compared to models trained on individual tasks separately (Li et al. (2023)).
One of the most common methods in MTL is hard parameter sharing, which involves sharing hidden layers between all tasks while keeping several task-specific output layers (Li et al. (2023)). It has been found effective in reducing model complexity and helping prevent overfitting, which is crucial for maintaining the fairness of predictions across different demographic groups as shown in Li et al. (2023) where different sub-populations are divided into separate tasks and the gradients of various prediction tasks during neural network back-propagation are dynamically modified to ensure fairness. This approach was tested on data from sepsis patients and was shown to significantly improve fairness disparity on the test data.
In the 2022 publication Gao et al. (2022), propose a novel approach to addressing gender bias in machine learning models through the framework of negative multi-task learning (NMTL). This technique is designed to actively reduce bias by learning tasks that negatively correlate with the biased outcomes, thereby mitigating undesirable bias propagation in predictive models. The concept of NMTL involves training models on tasks that are inversely related to biased decision-making processes, effectively using these negative correlations to counteract bias. The study presents results that demonstrate significant improvements in fairness metrics in the context of gender bias across various datasets.
Some studies have also integrated fairness directly into the MTL framework by treating fairness as an auxiliary task. For example, Oneto et al. (2019) proposes a framework where fairness is incorporated as an auxiliary objective alongside the primary classification tasks. This allows the model to leverage shared information across tasks to reduce bias inherently present in single-task learning environments. The authors showed that their approach improved fairness metrics without a substantial loss in accuracy through experiments on synthetic and real-world datasets.
While there are works thatlay a significant foundation in using MTL for fair classification (Oneto et al. (2019); Gao et al. (2022); Li et al. (2023)), our proposed MTL approach introduces several key improvements by integrating uncertainty estimation techniques, specifically MC Dropout, to further refine the understanding of how different input features influence the model’s predictions. This approach not only helps in identifying bias but also in dynamically adjusting the model’s reliance on various features based on their contribution to unfairness. We also extend this concept of MTL by incorporating Pareto optimality principles, which allows our model to explore a broader spectrum of potential solutions that balance multiple fairness metrics and accuracy more dynamically. This is particularly useful in complex real-world scenarios where different fairness considerations might conflict. This ensures that our proposed approach is suitable for a wider range of applications due to this adaptability feature.
2.3 Model Uncertainty for Bias Mitigation
The utilization of model uncertainty to mitigate biases in machine learning has been explored in several recent studies, highlighting different approaches.
In their work, Stone et al. (2022) introduce a method that incorporates epistemic uncertainty which refers to uncertainty caused by a lack of knowledge of the loss function used for training visual models (Hüllermeier and Waegeman (2021)). Their approach uses uncertainty measurements to weigh the loss function, prioritizing learning from less certain instances which are hypothesized to be less biased. This method helps reduce bias in visual recognition tasks by modifying how the model learns from data points with high uncertainty (Stone et al. (2022)). Our approach borrows from this idea by utilizing MC Dropout to determine model uncertainty in our prediction models.
Another approach is also introduced by Heuss et al. (2023) where model uncertainty is applied to mitigate bias in ranking algorithms. They propose a framework where predictive uncertainty is used to adjust the rankings to be fairer across different groups, ensuring that the uncertainty in model predictions does not disproportionately affect any single group (Heuss et al. (2023)). Almeida et al. (2021) on the other hand argued that uncertainty in class boundaries for classification tasks can lead to biased and inconsistent decisions. Their method reduces this specific type of uncertainty to improve both the fairness and the accuracy of the model.
Based on the evidence from prior works (Stone et al. (2022); Heuss et al. (2023); Almeida et al. (2021)), our research integrates uncertainty into our proposed MTL framework to enhance bias mitigation across multiple domains.
3 Fairness Terminology